A String is stretched and fastened to two points l apart. Motion is started by
displacing the string in the form
1
Expert's answer
2021-09-20T01:47:25-0400

"The \ displacement \ y(x, t) \ of \ the \ point \ of \ the \ string \ at \ a \\\ distance \ x \ from \ the \ left \ end \ 0 \ at \ time \ t \ is \ given\\ \ by \ the \ equation 
 
\frac{\partial^{2} y}{\partial t^{2}}=a^{2} \frac{\partial^{2} y}{\partial x^{2}} \\
 
Since \ the \ ends \ of \ the \ string \\ \ x=0 \ and \ x=l \ are \ fixed, \ they \ do \ not \ undergo\\ \ any \ displacement \ at \ any \ time.\\

Therefore, y(0, t)=0 \ for \ t \geq 0 \ and \ y(l, t)=0 \ for \ t \geq 0 \\
Since \ the \ string \ is \ released \ from \ rest \ initially, \\\ that \ is, \ at \ t= 0 , the \ initial \ velocity \ of \ every \ point \ of \ the\\ \ string \ in \ the \ y - direction \ is \ zero.\\
Hence, \ \frac{\partial y}{\partial t}(x, 0)=0 , for \ 0 \leq x \leq l\\
 
\frac{\partial^{2} y}{\partial t^{2}}=a^{2} \frac{\partial^{2} y}{\partial x^{2}}\\
 
Solving \ the \ PDE \ by \ Separation \ of \ Variables,\\"


"Let y=X(x) T(t) 
 
 \frac{\partial y}{\partial x}=X^{\prime}(x) T(t) \\
 \frac{\partial^{2} y}{\partial x^{2}}=X^{\prime \prime}(x) T(t) \\
 \frac{\partial y}{\partial t}=X(x) T^{\prime}(t) \\
 \frac{\partial^{2} y}{\partial t^{2}}=X(x) T^{\prime \prime}(t) \\
 X(x) T^{\prime \prime}(t)=a^{2} X^{\prime \prime}(x) T(t) \\
 \frac{X^{\prime \prime}(x)}{X(x)}=\frac{T^{\prime \prime}(t)}{a^{2} T(t)}=K^{2} \\
 \frac{X^{\prime \prime}(x)}{X(x)}=K^{2} \\
 X^{\prime \prime}(x)-K^{2} X(x)=0\\
 
Solution \,of \,X(x) \, is\,
 
 X(x)=C_{1} \cos (K x)+C_{2} \sin (K x) \\
 \frac{T^{\prime \prime}(t)}{a^{2} T(t)}=K^{2} \\
 T^{\prime \prime}(t)-a^{2} K^{2} T(t)=0"


"T(t)=C_{3} \cos (a K t)+C_{4} \sin (a K t)\\
 
Thus, \ y(x, t)=\left(C_{1} \cos (K x)+C_{2} \sin (K x)\right)\left(C_{3} \cos (a K t)+C_{4} \sin (a K t)\right) \\
By \ initial \ condition, \ \frac{\partial y}{\partial t}(x, 0)=0\\ 
 
 \frac{\partial y}{\partial t}=\left(C_{1} \cos (K x)+C_{2} \sin (K x)\right)\left(-C_{3} a K \sin (a K t)+C_{4} a K \cos (a K t)\right) \\
 \frac{\partial y}{\partial t}(x, 0)=C_{4} a K\left(C_{1} \cos (K x)+C_{2} \sin (K x)\right)=0 \\
 \Longrightarrow C_{4}=0\\
 
By \ the \ initial \ condition \ y(0, t)=0 \\ 
 
 y(x, t)=\left(C_{1} \cos (K x)+C_{2} \sin (K x)\right)\left(C_{3} \cos (a K t)+C_{4} \sin (a K t)\right)=0 \\
 y(0, t)=C_{1}\left(C_{3} \cos (a K t)+C_{4} \sin (a K t)\right)=0 \\
 \Longrightarrow C_{1}=0 \\
 \therefore y(x, t)=C_{2} \sin (K x) C_{3} \cos (a K t)\\
 
By \ the \ initial \ condition \ y(l, t)=0 \\
 
 y(l, t)=C_{2} \sin (K l) C_{3} \cos (a K t)=0 \\
 \Longrightarrow \sin (K l)=0 \\
 \sin (K l)=\sin (n \pi) \quad \text { for } n \geq 1\\"


"K=\frac{n \pi}{l} \\
 \therefore y(x, t)=C_{2} C_{3} \sin \left(\frac{a n \pi x}{l}\right) \cos \left(\frac{a n \pi t}{l}\right)=\lambda \sin \left(\frac{a n \pi x}{l}\right) \cos \left(\frac{a n \pi t}{l}\right)\\
 
The \ general \ solution \ is \ thus,\\ 
 
y(x, t)=\sum_{n=1}^{\infty} \lambda_{n} \sin \left(\frac{a n \pi x}{l}\right) \cos \left(\frac{a n \pi t}{l}\right)
 
Where, \ \lambda_{n} \ can \ be \ related \ to \ the \ Fourier \\
Coefficients \ by \ incorporating \ the \ initial \ conditions.\\
 
y(x, 0)=\sum_{n=1}^{\infty} \lambda_{n} \sin \left(\frac{a n \pi x}{l}\right) \cos \left(\frac{a n \pi \times 0}{l}\right)=\sum_{n=1}^{\infty} \lambda_{n} \sin \left(\frac{a n \pi x}{l}\right) \\
 
By \ Fourier's \ Half-range \ sine \ series, \ we \ derive \ \\ 
 
\lambda_{n}=\frac{2 a}{l} \int_{0}^{\frac{1}{a}} y(x, 0) \sin \left(\frac{a n \pi x}{l}\right) \mathrm{d} x"